FTIR OBSERVATION AND DFT STUDY OF THE AlC_{3} and AlC_{3}Al LINEAR CHAINS TRAPPED IN SOLID Ar

\underline{SARAH A. BATES}, C. M. L. RITTBY, and W. R. M. GRAHAM, Department of Physics and Astronomy, Texas Christian University, Fort Worth, TX 76129.

The vibrational spectra of linear AlC_{3} and AlC_{3}Al were observed after trapping the products of the dual laser evaporation of aluminum and carbon rods in solid Ar at ∼10 K. Fourier transform infrared (FTIR) measurements of ^{13}C isotopic shifts are in good agreement with the predictions of density functional theory (DFT) B3LYP/6-311G+(3df ) calculations and have enabled the first identification of the ν_{3}(σ_{u})=1624.0 and ν_{4}(σ_{u})=528.3 cm^{−1} fundamentals of linear AlC_{3}Al and the tentative assignment of the ν_{2}(σ)=1210.9 cm^{−1} fundamental of linear AlC_{3}.